OpenStudy (anonymous):

Two angles in a triangle measure 40 degrees and 55 degrees . What is the measure of the third angle ?

OpenStudy (anonymous):

85

OpenStudy (anonymous):

Just explain the procedure and steps of how to do this

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

sum of angle in a triangle is 180 so let the unknown angle be x so we have 40+55+x=180 95+x=180 x=180-95 x=85

OpenStudy (anonymous):

Thanks